A person is lying on a diving board 3.00 m above the surface of the water in a swimming pool. She looks at a penny that is on the bottom of the pool directly below her. To her, the penny appears to be a distance of 7.00 m from her. What is the depth of the water at this point?



1
Expert's answer
2021-11-24T12:02:09-0500

"n=1.33,"

"\Delta h=H-h=7-3=4~m,"

"h'=n\Delta h=5.3~m."